# Extreme Precipitation Nowcasting using Multi-Task Latent Diffusion Models

### v2

@@ Abstract
Deep learning models have made remarkable strides in precipitation prediction, yet they continue to struggle with capturing the spatial details of the features of radar images, particularly over high precipitation intensity areas. This shortcoming is evident in the form of low forecast accuracy in the spatial positioning of radar echo images across varying precipitation intensity regions. To address this challenge, we introduce the multi-task latent diffusion model (MTLDM), a novel approach for precipitation prediction. The basic concept of the MTLDM is based on the understanding that the radar image representing precipitation is the result of multiple factors. Therefore, we adopt a divide-and-conquer approach, that is, we decompose the radar image using decomposition technology and then predict the decomposed sub-images separately. We conceptualize the precipitation image as a composition of various components corresponding to different precipitation intensities. The MTLDM decomposes the precipitation image into these distinct components and employs a dedicated task to predict each one. This method enables spatiotemporally consistent prediction of real-world precipitation areas up to 5–80 min in advance, outperforming existing state-of-the-art techniques across multiple evaluation metrics.
@@
Precipitation Prediction,Multi-task Prediction,Deep Learning,Radar Image Decomposition,Latent Diffusion Model.
@@ I Introduction
Short-term precipitation is difficult to forecast accurately due to its rapid formation and dissipation. Short-term forecasts, i.e., 1–2 h in advance, can capture the position of weather systems and predict the range of precipitation in a timely manner, which is crucial for promptly informing the public about the time, location, and intensity of rainfall[1, 2]. However, as global climate change and urbanization accelerate, precipitation patterns are constantly changing, which poses higher demands on precipitation prediction technology. Deep learning models[3, 4, 5] have been shown to have a more powerful predictive performance than numerical weather prediction (NWP)[6]. However, as mentioned above, deep learning models still have serious deficiencies in the accuracy of spatiotemporal information prediction, making it difficult for them to meet practical needs. Therefore, researching high-precision forecasting models not only has important scientific value but also has broad application prospects. The inaccuracy of spatiotemporal information prediction is mainly manifested in the blurring of radar echo images generated during the prediction process, which is the result of inadequate convergence during the training process due to imperfect design structures of deep learning models. Moreover, the poor predictive performance of the details of spatial features is reflected in the low critical success index (CSI) of the generated radar echo images. Currently, deep learning models can hardly exceed a CSI of 0.5 under heavy precipitation conditions ($\geq$16 mm/h)[3, 7]. This situation indicates that deep learning models have significant problems in predicting the detailed features of images and cannot accurately predict precipitation conditions. As the prediction duration increases, the spatial prediction accuracy decreases dramatically. The inaccuracy of long sequence prediction is one of the main problems faced in all current deep learning research. Existing studies and our experiments clearly show that when the prediction duration exceeds 1 h, the CSI indicators for both heavy and light precipitation decrease sharply. Currently, the mainstream generative backbone network is the encoder–decoder (E–D) structure. We can use E–D networks with temporal capture capabilities (UNet and transformer networks) for precipitation prediction. Google uses UNet to predict precipitation[1], but the generated image quality easily suffers from blurring or loss of detail (such as the disappearance of small islands on the map). To improve the performance of spatiotemporal information prediction, Trebing et al.[8] introduced the Self-Attention(SA) into UNet and reduced the number of parameters in convolution-based models. In 2021, Ravuri Name et al. [3] achieved breakthrough performance using generative adversarial networks (GANs) to train a Convolutional Gated Recurrent unit (ConvGRU)-based E–D model (called DGMR) for precipitation prediction. This research attracted widespread attention to the use of deep networks in the field of precipitation prediction (i.e., weather prediction). Ling et al. [9] proposed a nested temporal UNet (TU2Net) that deepens the network hierarchy, significantly improving the quality of the predicted spatiotemporal information compared to the DGMR. Since deep networks are data-driven learning methods and cannot adhere to inherent physical laws, Zhang et al. [7] proposed NowcastNet, a nonlinear nowcasting model for extreme precipitation, which unifies physical evolution schemes and conditional learning methods into an end-to-end E–D network. However, like DGMR, NowcastNet still has low CSI values for heavy rainfall, and the CSI value for heavy precipitation($\geq$64 mm/h) decreases sharply to around 0.1 in the first 30 min.
@@
The diffusion probabilistic model (DPM) is a generative model that has emerged in the past two years and has been widely applied in video generation, image segmentation, knowledge reasoning, and other applications [10, 11]. It has been proven to outperform GAN models in various application areas. However, literature on applying the DPM to precipitation prediction only began to appear at the end of 2023. Zhihan et al. [12] proposed using the DPM for precipitation prediction, and its CSI value outperforms that of the GAN-based DGMR (both methods use UNet as the backbone network). Zhao et al.[13] combined the transformer with the DPM for precipitation prediction, confirming that the transformer achieves better CSI values than UNet. However, as the DPM has just begun to be applied to weather prediction such as rainfall, there is still significant room for improvement in its predictive performance. For instance, experimental data [13] show that when the precipitation is $>$ 3.5 mm/h, the CSI values of all of the methods become very low in heavy precipitation areas, and their average values rarely exceed 0.3.
@@
Figure 1: Radar threshold decomposition and feature extraction. (a) Original radar image, and (b and c) radar sub-images.
In radar reflectivity precipitation images, different colors or brightness values usually represent different precipitation intensities. Redder areas may indicate a higher precipitation intensity, while bluer areas indicate a lower precipitation intensity or no precipitation. We threshold-decompose the radar image into sub-images with different amounts of rainfall, and each sub-image represents a specific component or feature in the image. When using deep networks for image feature extraction (or calculating model loss), the pixel features of local areas (or the entire) in radar images are computed. When the deep network acts on the features of the radar sub-images, it masks out the interference from elements of other radar sub-images, thus making the features extracted by the model more accurate and also making the network easier to train. This method helps reduce noise and interference, allowing the model to focus more on tasks of different precipitation intensity levels. Fig.1 presents a schematic diagram of the radar threshold decomposition and feature extraction.
@@
To improve the spatiotemporal accuracy of precipitation prediction, we adopted a hierarchical prediction method that decomposes complex problems, thus breaking down the precipitation image into different components. Fig.2 shows overview of the proposed multi-task latent diffusion model (MTLDM) for precipitation prediction. After the input radar image sequence is predicted by the latent space diffusion model, different decoders are used to generate different precipitation sub-images from the predicted latent variables. Simultaneously, we also predict and generate the overall image. Based on the radar precipitation intensity, we set three thresholds ($Th_{1},Th_{2},andTh_{3}$) to segment a single radar image frame into different precipitation sub-images to train the Denoising Net and use a multi-task structure to predict the radar images of different precipitation intensities.
@@ III-B Model details
Fig. 3 shows the details of the MTLDM precipitation forecast model, which shares the same backbone network with the spacetime separable latent diffusion model (SSLDM, which is our previous work)[25]. The condition consists of four frames of radar images, which are used to predict 16 frames of radar images. The diffusion model adopts the latent diffusion model [26] meaning that the diffusion prediction encodes the input components (referred to as the conditions) and performs prediction in the latent space and then finally restores the predicted radar images through a decoder. The conditions are encoded through multiple encoders, forming multi-perspective features, which are then concatenated with Gaussian noise and fed into the Denoising NET. In the latent space encoding, the intensity structure information about the radar images is considered and is used to highlight the high-frequency information of the image and to enhance the prediction accuracy for areas of heavy rainfall.
@@
Multi-Task Learning: The model seems to be trained on multiple tasks simultaneously. This could involve predicting different aspects of the precipitation intensity.
AutoEncoder We use the Autoencoder used by Rombach et al.[26] and the intensity structural information of the radar to encode the latent space variables [25]. This module uses a ResnetBlock stack based on a convolutional network, and the KL-enalty factor is used during the training. To improve the reconstructing quality of the encoder, we added the structural similarity index similarity (SSIM[27] and learned perceptual image patch similarity (LPIPS) [28] losses to train the network. Unlike traditional methods, the SSIM evaluates the image brightness and contrast, as well as the structural information. Using the SSIM and LPIPS as loss functions, high-frequency signals in extreme precipitation events can be captured and reconstructed better, maximizing the performance of the latent of the DM. Thus, the AutoEncoder loss can be expressed as follows:
@@
III-C Denoising NET and Multiview Condition NET
As the main component of the diffusion model, Denoising NET receives the noisy latent space representation as the input, gradually denoises it through the network, and finally generates a clean latent space representation. The Denoising NET consists of a spatiotemporal separation structure of multiple temporal self-attention residual blocks (TSRBs) + spatial attention + temporal attention modules (Fig. 4(a)). The self-attention mechanism[29] has become a core component of state-of-the-art models [30] due to its powerful and flexible modeling capabilities. Spatial attention focuses on the spatial dependencies within the latent space representation and models remote spatial context information through the self-attention mechanism. Temporal attention[31, 32] regards the input as a time series and captures the temporal dependency of noise evolution by modeling the noise level dimension through the self-attention mechanism. By applying self-attention along the time axis, the model can better capture the temporal information and dependencies in the input data. This mechanism allows the model to refer to the input states and features from previous time steps while processing the current input, enabling it to learn richer temporal feature representations from long sequences. Serial connections further enhance the spatiotemporal modeling capabilities of the denoising network, and stacking multiple TSRBs expands the network’s receptive field and feature representation capabilities.
Fig. 4(b) shows the detailed framework of the multiview condition NET. It takes a conditional radar frame with size $4\times 1\times 256\times 256$ as the input. The multiview condition NET consists of two branches . The left branch obtains the latent space condition through the latent space encoder($E_{2}$ ), and its structural parameters remain unchanged during the diffusion model training process. The right branch uses the pixel encoder to encode the original image space($E_{3}$). The main part of the network uses the spatial attention mechanism and the down-sampling layer to reduce the spatial scale of the condition to $1\times 12\times 32\times 32$ to obtain the pixel space condition.
@@ III-C Denoising NET and Multiview Condition NET
Finally, we fuse the spatiotemporal fused latent space condition and pixel space condition with the latent space noise using a concatenation operation as the input of the Denoising NET to guide the denoising process. The latent space condition encodes the high-level semantic and structural information of the input data on the latent space manifold, which can guide the denoising network to grasp the global distribution and evolution trend of the data at the macro level. The pixel space condition captures the local details and texture features of the input data in the original image space, which helps the denoising network to restore the fine structure and high-frequency details of the image at the micro-scale. The pixel space network participates in the end-to-end training process of the entire diffusion model, which can adaptively remove redundant features and obtain more effective features.
Temporal Shift Residual Block The TSRB (Fig.5) integrates a temporal shift operation[20] (the temporal shift module (TSM) achieves temporal modeling by exchanging information between adjacent frames by shifting some channels in the feature map along the temporal dimension. This method is based on 2-D convolution and can achieve a performance comparable to those of 3-D convolutional neural networks without additional computation while maintaining the computational complexity of 2-D convolutional neural networks), residual connection, and time embedding to form an efficient spatiotemporal feature extractor. It uses dual temporal shift operations, which apply a temporal shift at the input and output to effectively inject timing information into feature maps without increasing the number of parameters and the computational complexity. This innovative design enables 2-D convolution to efficiently process 3-D data, playing an important role in the spatiotemporal capture of denoising models.
Figure 5: TemporalShift ResBlock (TSRB)
@@ IV-A Datasets and evaluation metrics
The experiment was conducted on two datasets. MRMS dataset: The second dataset comes from the multi-radar multi-sensor (MRMS) dataset for the United States[3]. We use the data for 2020–2021 as the training set and the data for 2022 as the testing set. Swedish dataset(SW)[25]: This dataset includes radar data from February 6, 2017, to November 12, 2021. We use the data from February 2017 to November 2020 as the training set, and the remaining data are used for validation and testing. This dataset is similar to grayscale images, with values ranging from 0 to 255, where 0 represents no echo, 255 represents no data (indicating a measurement point within radar coverage but with no reported echo), and the areas outside the radar coverage. For both of these datasets, the images used are cropped from the radar stream to a size of $256\times 256$ pixels.
@@ IV-B Baselines
We compared several models, including the Python framework for short-term ensemble prediction system (PySTEPS) (a popular machine learning method for precipitation prediction) [33], the DGMR model (a well-known conditional GAN for short-term precipitation prediction) and the timed racing diffusion model (TRDM) (a two-stage diffusion model-based method for precipitation prediction) [24]. The TRDM includes two versions: one that uses a super-resolution diffusion model for prediction in the second stage (called TRDMv1), and another that uses LDM for prediction (called TRDMv2). Two important metrics for evaluating precipitation are used: the CSI [34, 3, 7] and the continuous ranked probability score (CRPS)[3, 35]. The CSI is a commonly used statistical indicator for evaluating the accuracy of categorical forecasts and is often used to assess the prediction accuracy of weather events such as precipitation, storms, and snowfall. In evaluation, a higher CSI value indicates a higher accuracy of categorical forecasts. In precipitation assessment, the CRPS helps evaluate the consistency between the predicted precipitation probability distribution and the actual observed rainfall. Its advantage is that it considers both the precision and uncertainty of the predictions, making it particularly suitable for evaluating the performances of meteorological models. A smaller CRPS value indicates a better consistency between the predicted probability distribution and the observed values, suggesting a higher prediction accuracy.
Figure 7: Comparison of CSI Across Prediction Intervals for Different Precipitation Intensity Thresholds on the MRMS dataset on the SW dataset
Figure 8: Comparison of CRPS Across Prediction Intervals for Different Precipitation Intensity Thresholds on the MRMS dataset on the SW dataset

### v3

@@ Abstract
Deep learning models have achieved remarkable progress in precipitation prediction. However, they still face significant challenges in accurately capturing spatial details of radar images, particularly in regions of high precipitation intensity. This limitation results in reduced spatial localization accuracy when predicting radar echo images across varying precipitation intensities. To address this challenge, we propose an innovative precipitation prediction approach termed the Multi-Task Latent Diffusion Model (MTLDM). The core idea of MTLDM lies in the recognition that precipitation radar images represent a combination of multiple components, each corresponding to different precipitation intensities. Thus, we adopt a divide-and-conquer strategy, decomposing radar images into several sub-images based on their precipitation intensities and individually modeling these components. During the prediction stage, MTLDM integrates these sub-image representations by utilizing a trained latent-space rainfall diffusion model, followed by decoding through a multi-task decoder to produce the final precipitation prediction. Experimental evaluations conducted on the MRMS dataset demonstrate that the proposed MTLDM method surpasses state-of-the-art techniques, achieving a Critical Success Index (CSI) improvement of 13-26%.
@@
Precipitation nowcasting, Extreme Precipitation, Diffusion model, Multi-task.
@@ I Introduction
Short-term precipitation is difficult to forecast accurately due to its rapid formation and dissipation. Short-term forecasts, i.e., 1–2 h in advance, can capture the position of weather systems and predict the range of precipitation in a timely manner, which is crucial for promptly informing the public about the time, location, and intensity of rainfall[1, 2]. However, as global climate change and urbanization accelerate, precipitation patterns are constantly changing, which poses higher demands on precipitation prediction technology. Deep learning models[3, 4, 5] have been shown to have a more powerful predictive performance than numerical weather prediction (NWP)[6]. However, as mentioned above, deep learning models still have serious deficiencies in the accuracy of spatiotemporal information prediction, making it difficult for them to meet practical needs. Therefore, researching high-precision forecasting models not only has important scientific value but also has broad application prospects. The inaccuracy of spatiotemporal information prediction is mainly manifested in the blurring of radar echo images generated during the prediction process, which is the result of inadequate convergence during the training process due to imperfect design structures of deep learning models. Moreover, the poor predictive performance of the details of spatial features is reflected in the low critical success index (CSI) of the generated radar echo images. Currently, deep learning models can hardly exceed a CSI of 0.5 under heavy precipitation conditions ($\geq$16 mm/h)[3, 7]. This situation indicates that deep learning models have significant problems in predicting the detailed features of images and cannot accurately predict precipitation conditions. As the prediction duration increases, the spatial prediction accuracy decreases dramatically. The inaccuracy of long sequence prediction is one of the main problems faced in all current deep learning research. Existing studies and our experiments clearly show that when the prediction duration exceeds 1 h, the CSI indicators for both heavy and light precipitation decrease sharply.
Currently, the mainstream generative backbone network is the encoder–decoder (E–D) structure, such as Trebing et al.[8] introduced the Self-Attention(SA) into UNet and reduced the number of parameters in convolution-based models. In 2021, Ravuri Name et al. [3] achieved breakthrough performance using generative adversarial networks (GANs) to train a Convolutional Gated Recurrent unit (ConvGRU)-based E–D model (called DGMR) for precipitation prediction. This research attracted widespread attention to the use of deep networks in the field of precipitation prediction (i.e., weather prediction). Ling et al. [9] proposed a nested temporal UNet (TU2Net) that deepens the network hierarchy, significantly improving the quality of the predicted spatiotemporal information compared to the DGMR. Since deep networks are data-driven learning methods and cannot adhere to inherent physical laws, Zhang et al. [7] proposed NowcastNet, a nonlinear nowcasting model for extreme precipitation, which unifies physical evolution schemes and conditional learning methods into an end-to-end E–D network. However, like DGMR, NowcastNet still has low CSI values for heavy rainfall, and the CSI value for heavy precipitation($\geq$64 mm/h) decreases sharply to around 0.1 in the first 30 min.
@@
The diffusion probabilistic model (DPM) is a generative model that has emerged in the past two years and has been widely applied in video generation, image segmentation, knowledge reasoning, and other applications [10, 11]. It has been proven to outperform GAN models in various application areas. However, literature on applying the DPM to precipitation prediction only began to appear at the end of 2023. Zhihan et al. [12] proposed using the DPM for precipitation prediction, and its CSI value outperforms that of the GAN-based DGMR (both methods use UNet as the backbone network). Zhao et al.[13] combined the transformer with the DPM for precipitation prediction, confirming that the transformer achieves better CSI values than UNet. However, as the DPM has just begun to be applied to weather prediction such as rainfall, there is still significant room for improvement in its predictive performance. For instance, experimental data [13] show that when the precipitation is $>$ 3.5 mm/h, the CSI values of all of the methods become very low in heavy precipitation areas, and their average values rarely exceed 0.3. The fundamental reasons why existing deep learning models perform poorly under heavy precipitation conditions can be attributed to two aspects: data characteristics and model structure. From a data perspective, heavy precipitation events occur less frequently in nature, resulting in a serious shortage of heavy precipitation samples in the training data[3, 7], resulting in an imbalanced data distribution problem. This imbalance causes the model to tend to optimize the prediction of common precipitation intensities during training, at the expense of accurately capturing rare heavy precipitation events. In addition, heavy precipitation usually exhibits high nonlinearity and uncertainty in time and space, making it particularly difficult to learn these complex patterns from limited samples. From the perspective of model structure, existing deep learning models with single-task architectures have difficulty in simultaneously processing the feature representation of precipitation of different intensities[14]. Light precipitation and heavy precipitation show significantly different patterns and evolution laws on radar echo maps, and a single network architecture has difficulty capturing these differentiated features at the same time. In addition, traditional models usually use global loss functions such as mean square error (MSE)[8, 15]. When faced with images with a small proportion of heavy precipitation, such functions tend to smooth the prediction results and cannot accurately reconstruct the spatial details and boundary information of high-intensity precipitation areas. This results in the CSI value in heavy precipitation areas being significantly lower than that in light precipitation areas, which seriously limits the practical value of the model in extreme weather warning.
@@
Figure 1: Radar threshold decomposition and feature extraction. (a) Original radar image, and (b and c) radar sub-images. (d) Decomposition of real radar image according to precipitation threshold (R represents precipitation)
To improve the spatiotemporal aceuracy of precipitation prediction, we adopteda hierarchical prediction method that decomposes complex problems, thus breaking down the precipitation image into different components. We use threshold decomposition technology to divide radar images into multiple sub-images representing different precipitation amounts, each of which corresponds to a specific component in the original image. When using deep neural networks to extract features, the network processes local or overall pixel features of radar images. By extracting features from each sub-image separately, the interference from other precipitation level images can be effectively shielded, thereby improving the accuracy of feature extraction and simplifying the model training process. This method helps reduce noise interference and allows the model to focus more on task learning at different precipitation intensity levels. Fig.1 presents a schematic diagram of threshold decomposition and feature extraction of radar images.
@@
Fig.2 shows overview of the proposed multi-task latent diffusion model (MTLDM) for precipitation prediction. After the input radar image sequence is predicted by the latent space diffusion model, different decoders are used to generate different precipitation sub-images from the predicted latent variables. Simultaneously, we also predict and generate the overall image. Based on the radar precipitation intensity, we set three thresholds ($Th_{1}$, $Th_{2}$, and $Th_{3}$) to segment a single radar image frame into different precipitation sub-images to train the Denoising Net and use a multi-task structure to predict the radar images of different precipitation intensities.
@@
In summary, the main contributions of the Multi-Task Latent Diffusion Model (MTLDM) proposed in this paper are as follows:
  • We proposed a ”divide and conquer” strategy based on decomposition technology, decomposing radar images into sub-images corresponding to different precipitation intensities according to rainfall intensity, and designed a dedicated decoding task for each sub-image, so that the decoder can more accurately capture the spatial characteristics of precipitation areas with different intensities.
  • Based on the latent diffusion framework, MTLDM directly performs the diffusion process in the compressed latent space and directly decodes the generated latent space results. For the trained prediction model, there is no need to additionally train the latent space diffusion prediction model, which significantly reduces the research cost.
  • The multi-task learning framework we designed can optimize multiple related objectives simultaneously, significantly improving the prediction performance in high-intensity precipitation areas ($\geqslant 8mm/h$) while maintaining the prediction accuracy in low-intensity precipitation areas.
@@ III-B Model details
Fig. 3 shows the details of the MTLDM precipitation forecast model, which shares the same backbone network with the spacetime separable latent diffusion model (SSLDM, which is our previous work)[15]. The condition consists of four frames of radar images, which are used to predict 16 frames of radar images. The diffusion model adopts the latent diffusion model [27] meaning that the diffusion prediction encodes the input components (referred to as the conditions) and performs prediction in the latent space and then finally restores the predicted radar images through a decoder. The conditions are encoded through multiple encoders called Multiview Condition Net, forming multi-perspective features, which are then concatenated with Gaussian noise and fed into the Denoising NET. In the latent space encoding, the intensity structure information about the radar images is considered and is used to highlight the high-frequency information of the image and to enhance the prediction accuracy for areas of heavy rainfall.
@@
Multi-Task Learning: In order to ensure that the diffusion prediction model predicts in the same latent space distribution and verify the performance advantage of the multi-task decoder, we fixed the parameters of the encoder and only updated the parameters of the decoder. The input of the multi-task decoder is the decomposed threshold image, and its training goal is to accurately reconstruct the image. Since the encoder parameters remain unchanged and the input condition of the MCN is still the undecomposed radar image, the trained latent space prediction model can run normally. For the predicted latent space information, we directly use the multi-task decoder to decode it to obtain the final result.
AutoEncoder:We use the Autoencoder used by Rombach et al.[27] and the intensity structural information of the radar to encode the latent space variables [15]. This module uses a ResnetBlock stack based on a convolutional network, and the KL-enalty factor is used during the training. To improve the reconstructing quality of the encoder, we added the structural similarity index similarity (SSIM)[28] and learned perceptual image patch similarity (LPIPS) [29] losses to train the network. Unlike traditional methods, the SSIM evaluates the image brightness and contrast, as well as the structural information. Using the SSIM and LPIPS as loss functions, high-frequency signals in extreme precipitation events can be captured and reconstructed better, maximizing the performance of the latent of the DM. Thus, the AutoEncoder loss can be expressed as follows:
@@
Multiview Condition NET: Fig. 4(b) shows the detailed framework of the multiview condition NET. It takes a conditional radar frame with size $4\times 1\times 256\times 256$ as the input. The multiview condition NET consists of two branches . The left branch obtains the latent space condition through the latent space encoder($E_{2}$ ), and its structural parameters remain unchanged during the diffusion model training process. The right branch uses the pixel encoder to encode the original image space($E_{3}$). The main part of the network uses the spatial attention mechanism and the down-sampling layer to reduce the spatial scale of the condition to $1\times 12\times 32\times 32$ to obtain the pixel space condition.
@@ IV-A Datasets and evaluation metrics
The experiments were conducted on two datasets. The multi-radar multi-sensor (MRMS) dataset for the United States[3] provided data from 2020–2021 for training and data from 2022 for testing. The Swedish dataset (SW)[15] includes radar data from February 6, 2017, to November 12, 2021, with data from February 2017 to November 2020 used for training, and the remaining data used for validation and testing. The SW dataset is formatted similarly to grayscale images, with values ranging from 0 to 255, where 0 represents no echo, and 255 represents no data (indicating either a measurement point within radar coverage but with no reported echo, or areas outside the radar coverage). For both datasets, we cropped the images from the radar stream to a size of 256×256 pixels.
@@ IV-B Baselines
We compared several models, including the Python framework for short-term ensemble prediction system (PySTEPS) (a popular machine learning method for precipitation prediction) [30], the DGMR model (a well-known conditional GAN for short-term precipitation prediction) and the timed racing diffusion model (TRDM) (a two-stage diffusion model-based method for precipitation prediction) [26]. The TRDM includes two versions: one that uses a super-resolution diffusion model for prediction in the second stage (called TRDMv1), and another that uses LDM for prediction (called TRDMv2).The parameter sizes of these models are presented in Table II. Two important metrics for evaluating precipitation are used: the CSI [31, 3, 7] and the continuous ranked probability score (CRPS)[3, 32]. The CSI is a commonly used statistical indicator for evaluating the accuracy of categorical forecasts and is often used to assess the prediction accuracy of weather events such as precipitation, storms, and snowfall. In evaluation, a higher CSI value indicates a higher accuracy of categorical forecasts. In precipitation assessment, the CRPS helps evaluate the consistency between the predicted precipitation probability distribution and the actual observed rainfall. Its advantage is that it considers both the precision and uncertainty of the predictions, making it particularly suitable for evaluating the performances of meteorological models. A smaller CRPS value indicates a better consistency between the predicted probability distribution and the observed values, suggesting a higher prediction accuracy.